OpenStudy (anonymous):

solve -1/2x> or equal to -1/4

OpenStudy (anonymous):

x >= 1/2

OpenStudy (anonymous):

So x belong to the interval [1/2 , inf )

OpenStudy (anonymous):

can you show me the steps I went brain dead :-)

OpenStudy (anonymous):

you will just multiply by -2 both sides then you will get x >= 1/2

OpenStudy (anonymous):

So x belong to the interval [1/2 , inf )

OpenStudy (anonymous):

Does that help you?

OpenStudy (anonymous):

@2girls1boy

OpenStudy (anonymous):

yes I am tring to complete my finale and I can't remember any more

OpenStudy (anonymous):

Do you need any help for any question?

OpenStudy (anonymous):

@2girls1boy

OpenStudy (anonymous):

yes I do have one

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

but you can close this question

OpenStudy (anonymous):

are you still here?

OpenStudy (anonymous):

yes solve by elimanation 5x+6y=43 7x-6y=-55, substitution 8x+9y=41 -9x+y=54

OpenStudy (anonymous):

the first one is two equations right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5x + 6y = 43 7x - 6y = -55 ------------- add two equations 12 x = -12 x = -1

OpenStudy (anonymous):

I thought I had to multiply thanks

OpenStudy (anonymous):

8x + 9y = 41 --> equation 1 -9x + y = 54 --> y = 54+9x in equation 1: 8x + 9(54+9x) = 41 8x + 486 + 81x = 41 89x = -445 x = -5

OpenStudy (anonymous):

You were right!

OpenStudy (anonymous):

hey girl got it?

OpenStudy (anonymous):

yes thanks again

OpenStudy (anonymous):

no problem dear Can you message me?

OpenStudy (anonymous):

@2girls1boy

OpenStudy (anonymous):

on the top of the page and left corner

OpenStudy (anonymous):

@2girls1boy

OpenStudy (anonymous):

found it @2girls1boy